# Complex Saddles of Charged-AdS Gravitational partition function

### v1

@@ III.2 Fixed charge: Canonical ensemble
where $S(r_{+})$ and $E(r_{+},q)$ are entropy and energy of the charge spacetime, given by the relations in eq.(54) respectively. In terms of $r_{+}$ variable (ignoring Jacobian), we get
@@
The dominant saddle configurations that extremizes the exponent $\tilde{I}_{CAdS}(r_{+})$ of the integral in eq.(73) are given by the saddle equation:
@@
$\frac{d}{(d-2)l^{2}}r_{+}^{2d-2}+r_{+}^{2d-4}=\frac{4\pi}{(d-2)\beta}r_{+}^{2d-3}+q^{2}>q^{2}\,\,\text{for}\,\,r_{+}>0.$ (76)
@@ III.2.1 $AdS_{4}$
For $|a|\geq 1$, the discriminant remains negative for all values of $b\neq 1$, (see figure 12), indicating the presence of one positive real saddle point, one negative real saddle point, and a pair of complex conjugate saddles. According to the Picard–Lefschetz method, the complex conjugate pair of saddles—corresponding to naked singular geometries— are irrelevant and do not contribute to the path integral, as they fail to satisfy the necessary condition, $\mathcal{H}(r_{+}^{s})=0$.The negative real saddle (also naked singularity)— and the positive saddle, both satisfy the necessary condition, $\mathcal{H}(r_{+}^{s})=0$. However, for this regime it doesnot lead to any stokes degeneracy (see figure 13) for $0<b<1$, but does for $1<b$. Similar to the earlier cases, we lift the stokes degeneracy by complexifying the Newton constant. After complexification lifting the degeneracy and noticing that the negative saddle lies outside the defining contour, as a consequence its dual/steepest ascent to the thimble does not intersect the original integration contour, it is also irrelevant and does not contribute to the path integral. Meanwhile, the positive real saddle satisfies the necessary condition and lies on the defining contour, making it a relevant saddle according to the Picard–Lefschetz prescription. Therefore, for $|a|\geq 1$ at all temperatures the partition function gets contribution from only the regular and no-naked singular geometry. This situation is being illustrated in the figures (13) and (14).
@@
and one negative real solution $r_{\rm neg}=-l/(3\sqrt{6})$. Applying PL analogous to the earlier case, we obtain that the negative saddle is irrelevant as it lies outside defining contour. Whereas the triply degenerate positive saddle satifies the necessary condition and lies on the defining contour as shown in the figure (16). Therefore, the triply degenerate positive black holes is relevant and contribute to the path integral. Similarly, along the curves $b=b_{+}$ or $b=b_{-}$ with $|a|<1$, the discriminant vanishes. However, in these cases, unlike the critical case, one finds a real negative saddle, a real positive saddle, and a doubly degenerate positive saddle. All of them satisfying the necessary condition $\mathcal{H}(r_{s}^{+})=0$, with $\mathfrak{h}(r_{+}^{s})\neq\mathfrak{h}(r_{+}^{s^{\prime}})$ for $s\neq s^{\prime}$, denoting the stokes degeneracy situation. Upon complexifying the Newton constant and applying the Picard–Lefschetz analysis, one finds that both distinct positive black hole saddles are relevant, whereas the negative horizon solution is irrelevant. The doubly degenerate saddle ($r_{d}$) in either the cases $b=b_{+}$ or $b=b_{-}$ is given by
@@
Figure 17: PL plot for the $|q|<q_{c}$ and $\beta_{-}<\beta<\beta_{+}$ case, where the discriminant $\Delta>0$. In this case, we have a negative saddle denoted by NS, while the three distinct positive saddles $r_{+}^{1},r_{+}^{2},r_{+}^{3}$ are represented by BH1, BH2 and BH3, respectively. The flows associated with the negative saddle and the positive saddle that suffer Stokes jump are represented using dashed lines, while flow lines associated with contributing saddles are denoted by solid lines. For this illustration we choose $a=1/2,b=0.8$, and $G=e^{i\epsilon}$ where $b_{-}\approx 0.6456$ and $b_{+}\approx 0.95325$.
For $b_{-}<b<b_{+}$, one can convince themselves from the plot that $\Delta>0$, implying that the equation admits four real saddle solutions—either one positive and three negative saddles, or three positive and one negative saddle, depending on the value $b$ takes. However, according to Descartes’ rule of signs, which states that in a regime where all the roots are real, it can have only one negative saddle and three positive saddles, one finds that the latter case occurs. Again we encounter the Stokes degeneracy, which is resolved by complexifying the Newton constant. Let $r_{+}^{-}$ denote the negative saddle and $r_{+}^{1}$, $r_{+}^{2}$, and $r_{+}^{3}$ denote the three positive solutions satisfying $r_{+}^{1}<r^{2}_{+}<r^{3}_{+}$ for $|a|<1$ and $b_{-}<b<b_{+}$ case. Employing the Picard–Lefschetz technique, we find that all positive saddles are relevant for either sign of $\epsilon$, see fig. (17). However, the intermediate saddle $r^{2}_{+}$ undergoes a Stokes jump and drops upon taking homology averaging as discussed in sec. (III.2.3).
@@
To summarize, we find that in various regimes of $(q,\beta)$, as mentioned in table. 3 complex/negative saddles describing naked singularity are always irrelevant by PL and don’t contribute to the partition function. Only real and positive solution contribute to the partition function.
@@ III.2.2 $AdS_{5}$ and $AdS_{6}$
In the limit $\beta\to\infty$ the odd term $r_{+}^{2d-3}$ in the saddle equation given by (75) can be dropped, that excites for an analytic understanding about the saddle geometries of $AdS_{5}$ and $AdS_{6}$ for arbitrary $q$. In this limit, the saddle equation reduces to
@@
Similar to the $AdS_{4}$, we proceed with the same approach of analyzing the discriminant of the above equation, which serves as a good candidate to identify the nature of the saddle point solutions (The nature of the saddles for arbitrary $\beta>0$ with various regimes of $q$ is illustrated in Fig.(18)) for nature of saddle for given with any value of $\beta>0$) . For the quartic equation (the second equation in 86), the discriminant $\Delta$ is found to be negative for any value of $q$. This implies that, in the extremely low temperature limit, i.e., $\beta\to\infty$, $AdS_{6}$ has one real positive saddle, a real negative saddle and the rest are complex conjugate pairs. Implementing the Picard Lefschetz technique, we obtain that only the real positive saddle is relevant and contributes to the partition function. However, unlike $AdS_{6}$ case, the sign of $AdS_{5}$ discriminant depends on the value of charge $q$. To be precise, the discriminant of the cubic equation $\Delta$ (first equation in 86) is given by:
@@
For $q^{2}<l^{2}/27$, the above discriminant is negative, while for $q^{2}>l^{2}/27$ it is positive and vanishes for $q^{2}=l^{2}/27$. Irrespective of the discriminant sign, there always exists one real positive saddle, one real negative saddle, and the rest are complex conjugate saddles. The nature of Among these, only the real positive saddle is the relevant and contributing saddle to the path integral. These situations are illustrated in the figure (19).
@@ III.2.3 Homology averaging of Lefschetz Thimbles
Similar to the earlier case, in this section we discuss about the Stokes jump and homology averaging for the canonical ensemble. We notice when one fixes the charge; for the range of parameters $|q|<q_{c}$ and $\beta_{-}<\beta<\beta_{+}$, there are four saddles, specifically: one negative ($r_{+}^{-}$ ) and three positive solutions ($r_{+}^{1}<r_{+}^{2}<r_{+}^{3}$). Picard-Lefschetz analysis reveals the existence of a Stokes line among these saddles. To break this Stokes degeneracy, we complexify $G=|G|e^{i\epsilon}$ which leads to the discontinuous jump in the thimbles for $\epsilon>0$ and $\epsilon<0$ as shown in the figure (17). The homology jump in thimbles associated with the saddles can be expressed as
$\left(\begin{array}[]{c}\mathcal{J}(r_{+}^{-})\\ \mathcal{J}(r_{+}^{1})\\ \mathcal{J}(r_{+}^{2})\\ \mathcal{J}(r_{+}^{3})\end{array}\right)\rightarrow\left(\begin{array}[]{c|ccc}1&0&1&0\\ \hline\cr 0&1&1&0\\ 0&0&1&0\\ 0&0&1&1\end{array}\right)\left(\begin{array}[]{c}\mathcal{J}(r_{+}^{-})\\ \mathcal{J}(r_{+}^{1})\\ \mathcal{J}(r_{+}^{2})\\ \mathcal{J}(r_{+}^{3})\end{array}\right),\quad\text{for}\,\quad\epsilon^{+}\rightarrow\epsilon^{-}\,.$ (88)
Since the negative saddle $r_{+}^{-}$ is irrelevant and doesn’t contribute to the integral cycle, we refrain to consider the lower $3\times 3$ matrix in eq. (88, which is exactly the same as mentioned in eq. (29). Also note that $\mathcal{J}(r_{+}^{2})$ remains unchanged across the jump, hence the analysis will proceed same as in the case of high temperature AdS Schwarzschild case. Therefore, the contour of integration $\mathbb{D}\in(0^{+},\infty)$ can be expressed homologically two different ways as
$\begin{split}\mathbb{\mathbb{D}}=&\mathcal{J}(r_{+}^{1})-\mathcal{J}(r_{+}^{2})+\mathcal{J}(r_{+}^{3})\hskip 8.53581pt\text{for $\epsilon>0$}\\ =&\mathcal{J}(r_{+}^{1})+\mathcal{J}(r_{+}^{2})+\mathcal{J}(r_{+}^{3})\hskip 8.53581pt\text{for $\epsilon<0$}.\end{split}$ (89)
Figure 20: Plot of the canonical partition function exponent for various saddle points in the case $q=1/2$, shown for different values of $b$ in the range $b_{-}<b<b_{+}$, where $b_{-}\approx 0.6456$ and $b_{+}\approx 0.95325$. The plot clearly demonstrates that the intermediate saddle $r_{+}^{2}$ is always subdominant compared to the other saddles.
It is important to note that there is no $1/2$ factor in front of $\mathcal{J}(r_{+}^{1})$ as we consider the entire thimble associated with $r_{+}^{1}$. This difference comes because of an essential singularity at $r_{+}=0$ in the action when $q\neq 0$ and the thimble ends there. As explained earlier, upon doing the homology average, the contribution from the $r_{+}^{2}$ saddle (sub-leading) drops out.
@@ III.2.4 Thermodynamic stability analysis
From the above discussion, it is clear that the specific heat associated with the negative real saddle point, which appears in various regimes, is negative, whereas the specific heat of the complex saddles is complex. Now, we proceed to examine the behaviour of specific at the positive real saddle solutions in various ways by studying the sign of the denominator of $C_{q}$.
Simply, by examining the nature of saddle solutions and with a little bit of algebra, one may determine the sign of specific heat at the positive saddle solution. For $|q|\geq q_{c}$, at any temperature except at the critical temperature and $|q|<q_{c}$ with either $\beta<\beta_{-}$ or $\beta>\beta_{+}$, one finds a positive real black hole that has positive specific heat, indicating thermodynamic stability and is the sole relevant contributing saddle to the partition function. For $|q|<q_{c}$ we have three positive black hole solutions with horizon radii $r^{1}_{+}$, $r^{2}_{+}$, and $r^{3}_{+}$, satisfying $r^{1}_{+}<r^{2}_{+}<r^{3}_{+}$, all of which are relevant. We notice that the specific heat of small $r^{1}_{+}$ and large $r^{3}_{+}$ black holes have positive heat capacities, while medium black holes $r_{2}^{+}$ acquire a negative specific heat, indicating thermodynamic instability Chamblin:1999tk ; Ladghami:2024wkv . Nevertheless, the contribution of $r^{2}_{+}$ to the partition function drops out upon performing the homology averaging, as explained in sec. III.2.3. Hence, we reach the same conclusion as obtained before for AdS-Schwarzschild and the charged black hole in the grand ensemble.
Along the curves $\beta=\beta_{-}$ or $\beta=\beta_{+}$, with $|q|<q_{c}$, the specific heat at the non-degenerate positive saddle is positive, whereas the specific heat diverges at the doubly degenerate saddle point. Similarly, at the critical case $|q|\leq q_{c}$ and $\beta=\beta_{c}$, at the triply degenerate positive saddle $C_{q}$ diverges. At these points where saddles merge/specific heat diverges, the structure of Lefschetz thimbles changes discontinuously.

### v2

@@ III.2 Fixed charge: Canonical ensemble
where $S(r_{+})$ and $E(r_{+},q)$ are entropy and energy mentioned in eq.(55) respectively. In terms of $r_{+}$ variable (ignoring Jacobian), we get
@@
The dominant saddle configurations that extremize the exponent $\tilde{I}_{CAdS}(r_{+})$ of the integral in eq. (74) are given by the saddle equation:
@@
$\frac{d}{(d-2)l^{2}}r_{+}^{2d-2}+r_{+}^{2d-4}=\frac{4\pi}{(d-2)\beta}r_{+}^{2d-3}+q^{2}>q^{2}\,\,\text{for}\,\,r_{+}>0$ (77)
@@ III.2.1 $AdS_{4}$
For $|a|\geq 1$, the discriminant remains negative for all values of $b\neq 1$, (see figure 12), indicating the presence of one positive real saddle point, one negative real saddle point, and a pair of complex conjugate saddles. According to the Picard–Lefschetz method, the complex conjugate pair of saddles—corresponding to naked singular geometries— are irrelevant and do not contribute to the partition function, as they fail to satisfy the necessary condition, $\mathcal{H}(r_{+}^{s})=0$. The negative real saddle (also naked singularity)— and the positive saddle, both satisfy the necessary condition, $\mathcal{H}(r_{+}^{s})=0$. However, in this regime, it does not lead to any Stokes degeneracy (see figure 13) for $0<b<1$, but does for $1<b$. Similar to the earlier cases, we lift the Stokes degeneracy by complexifying the Newton constant. Lifting the degeneracy, we find that the negative saddle is irrelevant and does not contribute to the partition function. Meanwhile, the positive real saddle satisfies the necessary condition and lies on the defining contour, making it a relevant saddle according to the Picard–Lefschetz prescription. Therefore, for $|a|\geq 1$ at all temperatures the partition function gets contributions from only the regular and no-naked singular geometry. This situation is being illustrated in the figures (13) and (14).
@@
and one negative real solution $r_{\rm neg}=-l/(3\sqrt{6})$. Applying PL analogous to the earlier case, we obtain that the negative saddle is irrelevant as it lies outside defining contour. Whereas the triply degenerate positive saddle satifies the necessary condition and lies on the defining contour as shown in the figure (16). Therefore, the triply degenerate positive black holes is relevant and contribute to the partition function. Similarly, along the curves $b=b_{+}$ or $b=b_{-}$ with $|a|<1$, the discriminant vanishes. However, in these cases, unlike the critical case, one finds a real negative saddle, a real positive saddle, and a doubly degenerate positive saddle. All of them satisfying the necessary condition $\mathcal{H}(r_{s}^{+})=0$, with $\mathfrak{h}(r_{+}^{s})\neq\mathfrak{h}(r_{+}^{s^{\prime}})$ for $s\neq s^{\prime}$, denoting the stokes degeneracy situation. Upon complexifying the Newton constant and applying the Picard–Lefschetz analysis, one finds that both distinct positive black hole saddles are relevant, whereas the negative horizon solution is irrelevant. The doubly degenerate saddle ($r_{d}$) in either the cases $b=b_{+}$ or $b=b_{-}$ is given by
@@
Figure 17: PL plot for the $|q|<q_{c}$ and $\beta_{-}<\beta<\beta_{+}$ case, where the discriminant $\Delta>0$. In this case, we have a negative saddle denoted by NS, while the three distinct positive saddles $r_{+,1},r_{+,2},r_{+,3}$ are represented by BH1, BH2 and BH3, respectively. The flows associated with the negative saddle and the positive saddle that suffer Stokes jump are represented using dashed lines, while flow lines associated with contributing saddles are denoted by solid lines. For this illustration we choose $a=1/2,b=0.8$, and $G=e^{i\epsilon}$ where $b_{-}\approx 0.6456$ and $b_{+}\approx 0.95325$.
For $b_{-}<b<b_{+}$, one can convince themselves from the plot that $\Delta>0$, implying that the equation admits four real saddle solutions—either one positive and three negative saddles, or three positive and one negative saddle, depending on the value $b$ takes. However, according to Descartes’ rule of signs, which states that in a regime where all the roots are real, it can have only one negative saddle and three positive saddles, one finds that the latter case occurs. Again we encounter the Stokes degeneracy, which is resolved by complexifying the Newton constant. Let $r_{+}^{-}$ denote the negative saddle and $r_{+,1}$, $r_{+,2}$, and $r_{+,3}$ denote the three positive solutions satisfying $r_{+,1}<r_{+,2}<r_{+,3}$ for $|a|<1$ and $b_{-}<b<b_{+}$ case. Employing the Picard–Lefschetz technique, we find that all positive saddles are relevant for either sign of $\epsilon$, see fig. (17). However, the intermediate saddle $r_{+,2}$ undergoes a Stokes jump and drops upon taking homology averaging as discussed in sec. (III.2.3).
@@
To summarize, we find that in various regimes of $(q,\beta)$, as mentioned in table. 3 complex/negative saddles describing naked singularity are always irrelevant by PL and don’t contribute to the partition function. Only real and positive solutions contribute to the partition function, with the single exception of the $r_{+,2}$ saddle, whose contribution vanishes upon homology averaging.
@@ III.2.2 $AdS_{5}$ and $AdS_{6}$
In the limit $\beta\to\infty$ the odd term $r_{+}^{2d-3}$ in the saddle eq. (76) can be dropped, that excites for an analytic understanding about the saddle geometries of $AdS_{5}$ and $AdS_{6}$ for arbitrary $q$. In this limit, the saddle equation reduces to
@@
Similar to the $AdS_{4}$, we proceed with the same approach of analyzing the discriminant of the above equation, which serves as a good candidate to identify the nature of the saddle point solutions (The nature of the saddles for arbitrary $\beta>0$ with various regimes of $q$ is illustrated in Fig.(18)). For the quartic equation (the second equation in 87), the discriminant $\Delta$ is found to be negative for any value of $q$. This implies that, in the extremely low temperature limit, i.e., $\beta\to\infty$, $AdS_{6}$ has one real positive saddle, a real negative saddle and the rest are complex conjugate pairs. Implementing the Picard Lefschetz technique, we obtain that only the real positive saddle is relevant and contributes to the partition function. However, unlike $AdS_{6}$ case, the sign of $AdS_{5}$ discriminant depends on the value of charge $q$. To be precise, the discriminant of the cubic equation $\Delta$ (first equation in 87) is given by:
@@
For $q^{2}<l^{2}/27$, the above discriminant is negative, while for $q^{2}>l^{2}/27$ it is positive and vanishes for $q^{2}=l^{2}/27$. Irrespective of the discriminant sign, there always exists one real positive saddle, one real negative saddle, and the rest are complex conjugate saddles. The nature of Among these, only the real positive saddle is the relevant and contributing saddle to the partition function. These situations are illustrated in the figure (19).
@@
For finite $\beta$ , one expects more real roots and positive roots to appear. In general, the saddle eq. (76) can have a maximum of three real and positive roots ($\tilde{r}_{+,1},\tilde{r}_{+,2},\tilde{r}_{+,3}$). Similar to the $AdS_{4}$, these real roots will lie at $\tilde{r}_{+,1}<\tilde{r}_{+,2}<\tilde{r}_{+,3}$. In an exactly similar argument, the saddle $\tilde{r}_{+,2}$ will be subdominant, and the Lefschetz thimble ($\mathcal{J}(\tilde{r}_{+,2})$) corresponding to that saddle will show the Stokes jump.
@@ III.2.3 Homology averaging of Lefschetz Thimbles
Similar to the earlier analysis of the AdS Schwarzschild black hole in sec. (II.2.1), in this section we discuss about the Stokes jump and homology averaging for the canonical ensemble. We notice when one fixes the charge; for the range of parameters $|q|<q_{c}$ and $\beta_{-}<\beta<\beta_{+}$, there are four saddles, specifically: one negative ($r_{+}^{-}$ ) and three positive solutions ($r_{+,1}<r_{+,2}<r_{+,3}$). Picard-Lefschetz analysis reveals the existence of a Stokes line among these saddles. To break this Stokes degeneracy, we complexify $G=|G|e^{i\epsilon}$ which leads to the discontinuous jump in the thimbles for $\epsilon>0$ and $\epsilon<0$ as shown in the figure (17). The homology jump in thimbles associated with the saddles can be expressed as
$\left(\begin{array}[]{c}\mathcal{J}(r_{+}^{-})\\ \mathcal{J}(r_{+,1})\\ \mathcal{J}(r_{+,2})\\ \mathcal{J}(r_{+,3})\end{array}\right)\rightarrow\left(\begin{array}[]{c|ccc}1&0&1&0\\ \hline\cr 0&1&1&0\\ 0&0&1&0\\ 0&0&1&1\end{array}\right)\left(\begin{array}[]{c}\mathcal{J}(r_{+}^{-})\\ \mathcal{J}(r_{+,1})\\ \mathcal{J}(r_{+,2})\\ \mathcal{J}(r_{+,3})\end{array}\right),\quad\text{for}\,\quad\epsilon^{+}\rightarrow\epsilon^{-}\,.$ (89)
Since the negative saddle $r_{+}^{-}$ is irrelevant and doesn’t contribute to the integral cycle, we consider only the lower $3\times 3$ matrix in eq. (89, which is exactly the same as obtained earlier in eq. (29). Hence, the analysis will proceed in the same way as in the case of the high-temperature AdS Schwarzschild case. Note that the thimble $\mathcal{J}(r_{+,2})$ remains unchanged across the jump, as it is subdominant compared to the other saddles (see figure 20). Therefore, the contour of integration $\mathbb{D}\in(0^{+},\infty)$ can be expressed homologically two different ways as
$\begin{split}\mathbb{\mathbb{D}}=&\mathcal{J}(r_{+,1})-\mathcal{J}(r_{+,2})+\mathcal{J}(r_{+,3})\hskip 8.53581pt\text{for $\epsilon>0$}\\ =&\mathcal{J}(r_{+,1})+\mathcal{J}(r_{+,2})+\mathcal{J}(r_{+,3})\hskip 8.53581pt\text{for $\epsilon<0$}.\end{split}$ (90)
Figure 20: Plot of the canonical partition function exponent for various saddle points in the case $q=1/2$, shown for different values of $b$ in the range $b_{-}<b<b_{+}$, where $b_{-}\approx 0.6456$ and $b_{+}\approx 0.95325$. The plot clearly demonstrates that the intermediate saddle $r_{+,2}$ (red curve) is always subdominant compared to the other saddles.
It is important to note that there is no $1/2$ factor in front of $\mathcal{J}(r_{+,1})$ as we consider the entire thimble associated with $r_{+,1}$. This difference comes because of an essential singularity at $r_{+}=0$ in the action when $q\neq 0$ and the thimble ends there. The intersection number for the intermediate black hole saddle ($r_{+,2}$) is different for the two signs of $\epsilon$. In an exactly similar argument presented in sec. (II.2.1), upon doing the homology averaging, the contribution from the $r_{+,2}$ saddle (sub-leading) drops out.
@@ III.2.4 Thermodynamic stability analysis
From the above discussion, it is clear that the specific heat associated with the negative real saddle point, which appears in various regimes, is negative, whereas the specific heat of the complex saddles is complex. Now, we proceed to examine the behaviour of specific heat at the positive real saddles in various regimes by studying the sign of the denominator of $C_{q}$.
Simply, by examining the nature of saddle solutions and with a little bit of algebra, one may determine the sign of specific heat at the positive saddle solution. For $|q|\geq q_{c}$, at any temperature except at the critical temperature and $|q|<q_{c}$ with either $\beta<\beta_{-}$ or $\beta>\beta_{+}$, one finds a positive real black hole that has positive specific heat, indicating thermodynamic stability and is the sole relevant contributing saddle to the partition function. For $|q|<q_{c}$ we have three positive black hole solutions with horizon radii $r_{+,1}$, $r_{+,2}$, and $r_{+,3}$, satisfying $r_{+,1}<r_{+,2}<r_{+,3}$, all of which are relevant. We notice that the specific heat of small $r_{+,1}$ and large $r_{+,3}$ black holes are positive, while intermediate black holes $r_{+,2}$ acquire a negative specific heat (see figure 21), indicating thermodynamic instability Chamblin:1999tk ; Ladghami:2024wkv . Nevertheless, the contribution of $r_{+,2}$ to the partition function drops out upon performing the homology averaging, as explained in sec. III.2.3. Hence, we reach the same conclusion as obtained before for AdS-Schwarzschild and the charged black hole in the grand ensemble.
Along the curves $\beta=\beta_{-}$ or $\beta=\beta_{+}$, with $|q|<q_{c}$, the specific heat at the non-degenerate positive saddle is positive, whereas the specific heat diverges at the doubly degenerate saddle point. Similarly, at the critical case $|q|=q_{c}$ and $\beta=\beta_{c}$, $C_{q}$ diverges at the triply degenerate positive saddle. At these points where saddles merge/specific heat diverges, the structure of Lefschetz thimbles changes discontinuously.
Figure 21: A plot of specific heat capacity at fixed charge of $AdS_{4}$ black holes for $|q|<q_{c}$ and $\beta_{-}<\beta<\beta_{+}$. For illustration we choose $l=1,G=1$ and $q=q_{c}/2$. As explained in the text, the subdominant saddle $r_{+,2}$ has negative specific heat.